# FakeScope: Large Multimodal Expert Model for Transparent AI-Generated Image Forensics

### v1

@@ VI-E Ablation Studies
To investigate the influence of different data components and scales on training FakeScope, we further conduct ablation studies on the influence of three aspects during tuning with mPLUG-Owl2 [37]: (1) scale of training data; (2) collaboration of FakeChain and visual instructions; and (3) participation of multimodal contextual information.
1) Effects of Training Data Scale: To effectively handle data-scarce scenarios is crucial for forensic models, yet it remains increasingly challenging due to the high cost and difficulty of acquiring high-quality training data. To evaluate the effectiveness of FakeInstruct dataset, we randomly sample four subsets—80%, 60%, 40%, and 20%—while preserving the original data composition, and analyze the impact on training outcomes. As illustrated in Fig. 8, the resulting FakeScope variants exhibit remarkable resilience to performance degradation as the available training data decreases, demonstrating the strong utility of FakeInstruct. Notably, the transparency-related performance remains a significant improvement over the baselines, even at the extreme 20% data proportion, and continues to strengthen as the training data size grows. This trend highlights the effectiveness of the multimodal forensic knowledge embedded in FakeInstruct, further demonstrating that the reasoning data can substantially enhance the forensic capabilities of large multimodal models.
2) Effects of Collaborative Training: During training, we combine different subsets to jointly optimize a unified forensic foundation model. To validate the effectiveness of this strategy, we compare it against task-specific training focused solely on transparent forensic capabilities. As indicated by the results listed in Table VII, the collaborative training scheme—integrating long-form reasoning with granular visual instructions—yields significant improvements in model performance. Although the various subsets in FakeInstruct originate from a shared data source, this strategy remains cost-efficient and enables FakeScope to generalize more effectively across diverse forensic tasks, demonstrating greater stability across a broader knowledge domain.
@@
Training Strategy Detect Interpret Reason Analyze
none (baseline) 58.22% 0.423 0.296 0.368
only FakeChain 83.27% 0.593 0.576 0.350
full FakeInstruct 99.68% 0.638 0.614 0.540
TABLE VIII: Comparison on forensic detection ability between full FakeInstruct, only absolute judgment, and only contextual data. The AP (%) on each dataset is reported. The best performance of each column is highlighted in bold.
@@
FakeBench [7] 68.37 94.06${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+25.96}}}$ 84.13${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+15.76}}}$ 99.94${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+31.57}}}$
AntiFakePrompt [71] 60.06 90.12${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+30.06}}}$ 74.84${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+14.78}}}$ 98.90${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+38.84}}}$
AIGCDetecBenchmark [72] 54.84 81.65${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+26.81}}}$ 60.73${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+5.89}}}$ 90.23${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+35.39}}}$
WildRF [73] 52.73 82.34${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+29.61}}}$ 62.54${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+9.81}}}$ 90.05${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+37.32}}}$
SynthWildX [4] 47.41 82.78${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+35.37}}}$ 52.59${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+5.18}}}$ 86.93${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+39.52}}}$
3) Effects of Multimodal Awareness: A defining feature that sets FakeScope apart from conventional binary forensic detectors is its integration of rich multimodal contextual information on visual trace evidence during training. While these data are aimed at enhancing transparency-related forensic capabilities, they might affect the performance on simple detection. To validate whether such information also improves pure forensic detection, we compare the performance of models trained with and without it. As compared in Table VIII, incorporating contextual information not only improves the model’s transparency but also enhances detection accuracy and generalization, even when using the same amount of image data. In contrast, omitting absolute authenticity labels from training results in inconsistent improvements. These findings suggest that multimodal forensic cues can strengthen the model’s task awareness and fine-grained visual comprehension, enabling it to more effectively attend to forensic-relevant patterns. However, the model’s detection performance still heavily relies on absolute authenticity judgments, as relying solely on descriptive cues may introduce ambiguity into the decision-making process for authenticity.
@@ VII Conclusions and Future Work
In this paper, we propose a multimodal paradigm for transparent AI-generated image forensics, versatile in detection, in-depth analysis, and extended discussions. Based on the advanced cross-modal capability of large multimodal models (LMMs), we collect and leverage rich image forensic knowledge to enhance their forensic awareness. By emulating human causal inference based on visual trace evidence, we introduce a human-machine collaborative strategy and construct the first large-scale dataset containing structured forensic reasoning data. This reasoning foundation is further expanded into a million-scale visual instruction dataset, which is internalized into LMMs to improve their transparent forensic abilities on AI-generated images. Extensive experiments demonstrate consistent effectiveness of this paradigm across LMMs with varied meta frameworks. The proposed expert forensic model can reliably handle multiple forensic tasks under a unified paradigm, and exhibits strong generalization and in-the-wild capabilities. Besides, we discovered that LMMs can effectively align with hard authenticity labels and demonstrate impressive generalization on estimating numerical probabilities, even without explicit supervision.
This work not only presents a viable pathway toward transparent AI-generated image forensics, but also offers insights into the development of trustworthy AI risk management within the broader multimodal landscape. Moreover, the proposed model could serve as the source of generating cross-modal forensic information, which may further support the development of more reliable forensic systems. Notwithstanding the use of advanced models such as GPT-4V to distill forensic knowledge, it may still represent an early-stage effort in the evolving trajectory of large multimodal models. While some uncertainties remain, this study nonetheless demonstrates the effectiveness of multimodal information and the feasibility of forensic transparency. We intend to further explore the abilities of newly emerged LMMs and contribute model update to community. Looking ahead, future efforts may focus on enhancing the model’s few-shot adaptability and dialectical reasoning capability for better robustness and explanatory power in more complex forensic contexts.

### v2

@@ VI-E Ablation Studies
We analyze how data composition and scale affect FakeScope training with mPLUG-Owl2 [78]. Specifically, we study: (1) instruction scale, i.e., the effect of dataset size; (2) instructional diversity, by comparing FakeChain-only training against enriched multimodal supervision from FakeInstruct; (3) multimodal awareness, via detection performance with and without multimodal context. Besides, we further validate the (4) effectiveness of the proposed token soft-scoring strategy.
1) Effects of Training Data Scale: Data scarcity is a core bottleneck for forensic modeling because high-quality annotations are expensive to obtain. To assess the data efficiency of FakeInstruct, we sample 80%, 60%, 40%, and 20% subsets while preserving the original composition. As shown in Fig. 8, FakeScope degrades gracefully as training data decreases, indicating strong robustness of FakeInstruct. Even at 20%, performance remains clearly above baselines, and improves steadily with larger subsets. This trend suggests that the gains are driven primarily by the forensic knowledge encoded in FakeInstruct rather than data volume alone.
2) Effects of Instructional Diversity: We jointly train a unified forensic foundation model on multiple subsets and compare this setting with task-specific training focused only on transparent forensics. As shown in Table VII, collaborative training that combines long-chain reasoning with fine-grained visual instruction consistently improves performance, which benefits cross-task generalization. Full-schedule training improves all capabilities while reducing cross-task interference, indicating complementary rather than competitive interactions among forensic knowledge types. In particular, both analysis and attribution benefit from fine-grained supervision, whereas training with FakeChain alone provides no comparable improvement. These results validate the effectiveness of our visual instruction strategy.
Figure 8: Performance on FakeBench [40] concerning training data scale. The x-axis indicates the proportion of FakeInstruct used for training, and y-axis shows corresponding performance indicator of each ability dimension.
@@
Training Strategy Detect Interpret Reason Analyze Attribute
none (baseline) 58.22% 0.423 0.296 0.368 65.97%
only FakeChain 83.27% 0.593 0.576 0.350 68.43%
full FakeInstruct 99.68% 0.638 0.614 0.540 78.50%
TABLE VIII: Comparison on forensic detection ability among full FakeInstruct, only absolute judgment, and only contextual data containing trace evidence instructions. AP (%) on each dataset is reported.
@@
[…]
3) Effects of Multimodal Awareness: A key distinction between FakeScope and conventional binary detectors is the use of multimodal evidence for visual trace analysis. Although this design is intended to improve transparent forensics, its contribution to basic detection is not obvious. We therefore compare models trained with and without trace-evidence instructions. As reported in Table VIII, adding multimodal context improves both detection accuracy and generalization, even under the same image budget. By contrast, removing absolute authenticity labels yields less stable gains. These results indicate that multimodal forensic cues enhance task awareness and fine-grained visual understanding, whereas robust detection still requires explicit authenticity supervision.
TABLE IX: Ablations of the proposed token soft-scoring strategy: (i) auxiliary classifier; (ii) verbalized confidence; (iii) synonym ensembles for <auth_token>; and (iv) argmax vs. softmax scoring. We report mAP (%) on each dataset.
Exp. Model Strategy Fake Bench Antifake Prompt AD Bench. WildRF Synth WildX
(i) FakeScope (mPLUG-Owl2) coupled (ours) 99.94 98.90 90.23 90.05 86.93
de-coupled 98.67 95.20 87.69 85.42 83.89
FakeScope (LLaVA-v1.5) coupled (ours) 99.39 94.53 99.15 74.48 82.77
de-coupled 96.78 95.99 96.45 60.80 79.76
(ii) FakeScope (mPLUG-Owl2) token-based (ours) 99.94 98.90 90.23 90.05 86.93
verbalizegreedy 66.08 60.43 50.87 52.36 45.42
verbalizet=0.5 66.79 61.66 52.80 52.39 45.40
verbalizet=1.0 65.32 58.75 51.43 50.28 43.27
FakeScope (LLaVA-v1.5) token-based (ours) 99.39 94.53 99.15 74.48 82.77
verbalizegreedy 68.89 62.58 52.66 54.90 49.87
verbalizet=0.5 68.72 62.05 51.43 54.95 49.26
verbalizet=1.0 66.77 61.02 51.35 52.73 46.58
[…]
fake+artificial+synthetic $\leftrightarrow$ real+genuine+authentic 97.78 97.29 90.30 89.03 86.81
[…]
fake+artificial+synthetic $\leftrightarrow$ real+genuine+authentic 99.43 92.88 99.02 72.65 83.21
(iv) FakeScope (mPLUG-Owl2) softmax (ours) 99.94 98.90 90.23 90.05 86.93
argmax 95.66 98.47 63.57 55.36 78.78
FakeScope (LLaVA-v1.5) softmax (ours) 99.39 94.53 99.15 74.48 82.77
argmax 93.55 95.64 85.55 52.92 77.42
[…]
@@ VII Conclusions and Future Work
This paper proposes a multimodal paradigm for transparent AI-generated image forensics that integrates detection, in-depth analysis, and extended discussions within a unified framework. To support this paradigm, we build FakeChain, a large-scale forensic reasoning dataset created through human-machine collaboration, and extend it into FakeInstruct, the first multimodal instruction-tuning dataset for image forensics with 2 million visual instructions. On this basis, we develop FakeScope, an expert large multimodal model that achieves superior performance in both closed-ended and open-ended forensic tasks, while showing strong generalization and robustness in the wild. Notably, FakeScope demonstrates zero-shot quantitative detection without explicit numerical supervision, suggesting an emergent capacity for probabilistic reasoning.
More broadly, this work offers a practical path toward transparent and trustworthy AI-generated image forensics. Beyond accurate detection, FakeScope produces cross-modal forensic information, supporting explainable analysis. Our results also underscore the value of human-in-the-loop design and structured prompting for reliable knowledge distillation in LMMs. Although current progress still depends on advanced models such as GPT-4V and important uncertainties remain, the findings confirm the value of multimodal information for forensic transparency. Looking ahead, we intend to explore the abilities of newly emerged LMMs and contribute model updates to the community. Given the diminishing returns implied by scaling laws, future efforts may focus on enhancing few-shot adaptability and dialectical reasoning to boost robustness in more complex, unseen forensic scenarios.